OpenStudy (anonymous):

a circle has the following equation 2(x-3)^2+2y^2=8. find the center (h,k) and radius

OpenStudy (anonymous):

Find the formula for h,k

OpenStudy (anonymous):

First divide both sides by 2

OpenStudy (anonymous):

You now have (x - 3)^2 + y^2 = 4

OpenStudy (anonymous):

Center is (h,k) where the equation of the circle is (x - h)^2 + (y-k)^2 = r^2, and r = radius of the circle

OpenStudy (anonymous):

So, what is the center and radius of the circle?

OpenStudy (anonymous):

formula for h,k is (x-h)^2+(y-k)^2=r2

OpenStudy (anonymous):

idk im stuck